rubycutie:

If the 78.0 kg astronaut were in a spacecraft 6R from the center of the earth, what would the astronaut's weight be on earth

Shadow:

Hello @rubycutie

aidan2001:

Acceleration due to gravity on the surface of radius R = GM/R^2 = 9.8 m/s^2 weight at surface of earth = mg = 78 x 9.8 = 764.4 N acceleration due to gravity at 6R = GM/(6R)^2 = GM/R^2 /36 =9.8/36 = 0.27 m/s^2 weight on spacecraft = 78 x 0.27 = 21.2 N

aidan2001:

His weight on Earth doesn't depend on where he is at the moment. Weight = (mass) x (gravity) On the Earth's surface, the acceleration of gravity is 9.8 m/s². So, on the surface, the astronaut weighs                        (78 kg) x (9.8 m/s²)  =  764.4 Newtons  (about 172 pounds) . No matter where he goes, his mass doesn't change. But his weight does. If he's  6R  from the enter of the Earth, then he's 5 times as far from it as when he's standing on the surface.  The acceleration of gravity out there is                    ( 1 / 5² )  =  1/25  =  0.04 times the acceleration of gravity on the surface. His weight out there is                   (78 kg) x (9.8/25  m/s²)  =  30.6 Newtons  (about  6.9 pounds)